ryan can do a job in 12 days. after he has worked for 4 days ,he is joined by sandy.they finished the job together in 2 more days .how long would it take sandy to do the whole job herself?
Answer by richwmiller(17219)

4/12+2/12+2/s=1
6/12+2/s=1
2/s=6/12
2/s=1/2
s=4 days for Sandy alone